OpenStudy (itz_sid):

Differential Equation:

OpenStudy (itz_sid):

\[y'+4y=3\sin(e^{4x})\]

OpenStudy (itz_sid):

\[I(x) = e^{\int\limits 4dx} = e^{4x}\] \[\int\limits [e^{4x}y]' = \int\limits 3e^{4x}\sin(e^{4x})dx\]

OpenStudy (itz_sid):

u = 4x du = 4 dx \[\frac{ 3 }{ 4 } \int\limits e^u \sin(e^u)du\]

OpenStudy (itz_sid):

And now im stuck :/

OpenStudy (eliesaab):

What is the derivative of cos(e^u) with respect to u?

OpenStudy (eliesaab):

\[ \frac {d}{du}\cos(e^u)= -e^u \sin(e^u) \] You should be able to continue now

OpenStudy (itz_sid):

Oh is it the chain rule? The derivative of the thing times the inside?\[\cos(e^u)\times(e^u)\]

OpenStudy (itz_sid):

oops i meant -sin... lol

OpenStudy (eliesaab):

Your last integral is easy now

OpenStudy (itz_sid):

Yea got it. Thanks!

OpenStudy (sshayer):

put \[e^{4x}=u,e^{4x}4 ~dx=du,\] \[\int\limits 3 e^{4x} \sin (e^{4x})dx=\frac{ 3 }{ 4 }\int\limits \sin u~du=-\frac{ 3 }{ 4 }\cos u\]